OpenStudy (mgriffin16):

4 times the sum of a number and 3 is 9 less than twice the number?

OpenStudy (mgriffin16):

I must write this in an equation , not solving it

OpenStudy (welshfella):

I'm looking at that again. I got confused...

OpenStudy (welshfella):

You have 4 times the sum of a number and 3 that is written as 4(x + 3)

OpenStudy (welshfella):

now the value of the above is 9 less than twice the number so 4(x + 3) = 2x - 9

OpenStudy (mgriffin16):

Write the equation of the following statement: 4 times the sum of a number and 3 is 9 less than twice the number. This is the whole question

OpenStudy (welshfella):

you have it in my last post.

OpenStudy (mgriffin16):

A. 4(b + 3) = 9 - 2b B. 4b + 3 = 2b - 9 C. 4(b + 3) = 2b - 9 D. 4(b - 3) = 9 - 2b it can only be one of those four